OpenStudy (anonymous):

Find the radius of a circle with the given circumfrence. 15pi ft. A. 30 pi ft. B. 30 ft. C. 7.5 pi ft. D. 7.5 ft.

OpenStudy (snuggielad):

first of al do you know how to find the radius...what is the formula?

OpenStudy (anonymous):

They don't tell me. However, the formula to find the circumfrence is: C=2*pi*radius

OpenStudy (snuggielad):

when it says 15 pi feet does it mean pie as in 3.14?

OpenStudy (anonymous):

Yes..

OpenStudy (snuggielad):

i got this guys ok then you will do the formula in reverse...solve for R not c...so first find out what 15*3.1428 is and that is "C" then work out the problem... like this ({15*3.1428}/3.1428)/2=r...

OpenStudy (snuggielad):

tell me your answer...dont just do it and leave...

OpenStudy (anonymous):

Hold up you wrote out pi completely .... They only want me to solve with pi as 3.14 not 3.1428

OpenStudy (snuggielad):

...lol i promise...its not written out completely...lol...but it is to the fourth number...just solve it that way...it gives a clearer answer

OpenStudy (anonymous):

I got 7.49

OpenStudy (anonymous):

circumference = pi * diameter =pi * 2* radius you have the circumference. you have pi, and you have 2. just solve for the radius.

OpenStudy (anonymous):

Okay so basically it's this 15pi=2*pi*r

OpenStudy (anonymous):

they didn't times it by pi. the circumference either 15pi or 47 ft. its just in a different form. this means in its pi form, you only need to divide it by 2. Which gives you the radius of 7.5 ft.

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

Yea it does

OpenStudy (dean.shyy):

Remember that the circle circumference is 2 * radius * Pi. So, given 15 * Pi, that equates to the circle circumference being 2 * Pi * 7.5. Now, given this information provided to you, can you determine what the answer would be? If you have not figured it out, 7.5 * Pi best matches the third choice from the information provided here.

OpenStudy (anonymous):

whoops dean is correct.

OpenStudy (anonymous):

thank you for clearing that up

OpenStudy (anonymous):

Haha it's alright. Thanks Dean (: I would give you best response but I'm on an iPhone right now and it won't let me. Soo I'll give you best response when I get on my computer (: